# DAASH: A Meta-Attack Framework for Synthesizing Effective and Stealthy Adversarial Examples

### v1

@@ Abstract
Numerous techniques have been proposed for generating adversarial examples in white-box settings under strict $\ell_{p}$-norm constraints. However, such norm-bounded examples often fail to align well with human perception, and only recently have a few methods begun specifically exploring perceptually aligned adversarial examples. Moreover, it remains unclear whether insights from $\ell_{p}$-constrained attacks can be effectively leveraged to improve perceptual efficacy. In this paper, we introduce DAASH, a fully differentiable meta-attack framework that generates effective and perceptually aligned adversarial examples by strategically composing existing $\ell_{p}$-based attack methods. DAASH operates in a multi-stage fashion: at each stage, it aggregates candidate adversarial examples from multiple base attacks using learned, adaptive weights and propagates the result to the next stage. A novel meta-loss function guides this process by jointly minimizing misclassification loss and perceptual distortion, enabling the framework to dynamically modulate the contribution of each base attack throughout the stages. We evaluate DAASH on adversarially trained models across CIFAR-10, CIFAR-100, and ImageNet. Despite relying solely on $\ell_{p}$-constrained based methods, DAASH significantly outperforms state-of-the-art perceptual attacks such as AdvAD—achieving higher attack success rates (e.g., 20.63% improvement) and superior visual quality, as measured by SSIM, LPIPS, and FID (improvements of $\approx 11$, 0.015, and 5.7, respectively). Furthermore, DAASH generalizes well to unseen defenses, making it a practical and strong baseline for evaluating robustness without requiring handcrafted adaptive attacks for each new defense.
@@ 1 Introduction
Adversarial examples against deep neural networks (DNNs), which add imperceptible perturbations to clean images to induce misclassification (targeted/untargeted) Goodfellow et al. (2015a), have attracted significant attention from the research community. Various attack methods, including FGSM Goodfellow et al. (2015b), PGD Madry et al. (2018b), and CW Carlini and Wagner (2017), as well as adaptive evaluation frameworks such as AutoAttack Croce et al. (2020), have been developed to benchmark both defended and undefended models. However, the majority of these approaches prioritize imperceptibility by constraining perturbations within a small $\ell_{p}$-norm ball, with $\ell_{2}$ and $\ell_{\infty}$ norms being the most common ones.
Yet, minimizing $\ell_{p}$-norm does not always yield high perceptual similarity to the human visual system Zhang et al. (2018); Laidlaw and Feizi (2020); Sharif et al. (2018), despite imperceptibility being a fundamental design goal of stealthy adversarial examples. This disconnect has motivated the adoption of perceptually-aligned similarity metrics such as Structural Similarity Index (SSIM) Wang et al. (2004), Learned Perceptual Image Patch Similarity (LPIPS) Zhang et al. (2018), and Fréchet Inception Distance (FID) Heusel et al. (2017) to better assess visual quality Li et al. (2024); Chen et al. (2024b). However, attacks optimized under these perceptual metrics often suffer from reduced success rates—particularly against robust, defended models (see Table 1)—highlighting a critical trade-off and leaving significant room for improvement.
Designing effective attacks against a wide range of models, especially hardened ones, remains non-trivial and often requires deep, manual effort—an observation consistently echoed in the literature on $\ell_{p}$-norm attacks Carlini and Wagner (2017); Athalye et al. (2018); Croce and Hein (2020b). This complexity intensifies when attempting to simultaneously preserve high perceptual fidelity, thus underscoring the need for systematic and automated strategies that can generalize across models while aligning better with human vision.
Figure 1: Comparative visualization of adversarial perturbations across multiple attack methods (including DAASH).
@@
Our key insight is that although individual $\ell_{p}$-norms (e.g., $\ell_{2}$, $\ell_{\infty}$) do not fully capture human perception, they each correlate with different perceptual characteristics (e.g., contrast, texture, edge disruption) Sharif et al. (2018); Laidlaw and Feizi (2020). Therefore, by combining attack strategies optimized under different norms in a principled manner, we may achieve not only improved attack success but also better perceptual quality, particularly when targeting defended models where individual attack methods may fall short. Specifically, we propose DAASH (Differential Adversarial Attack SearcH)—a fully differentiable, multi-stage framework that learns to compose multiple $\ell_{p}$-norm–constrained base attacks using soft attention. DAASH operates in two stages: (1) it computes a soft attention distribution over adversarial examples generated from individual candidate attacks (e.g., PGD, CW) to weight and compose the final adversarial example; (2) it iteratively refines the perturbations across multiple stages by feeding the output from one stage as the input to the next. This iterative chaining is designed to escape poor local minima inherent to the non-convex landscape of adversarial optimization.
The attention weights are optimized end-to-end through a meta-loss that jointly considers (i) attack success (via minimizing the confidence score of the ground-truth class) and (ii) perceptual similarity (via maximizing SSIM). This allows the framework to dynamically adjust the influence of each candidate attack based on the target model. Unlike existing methods that are rigid in balance between imperceptibility (from human perception) and attack success, DAASH learns to trade off these objectives adaptively across different models and datasets.
Empirically, DAASH achieves a significant improvement in both attack success rate (ASR) and perceptual fidelity compared to the current state-of-the-art. On robust models trained on CIFAR-100 such as Cui et al. (2024), DAASH reaches an average ASR of 99.97% with an SSIM score of 94.43, significantly outperforming state-of-the-art perceptual attacks such as AdvAD Li et al. (2024) (ASR: 79.14%, SSIM: 83.18) and DiffAttack Chen et al. (2024b) (ASR: 77.24%, SSIM: 91.16). Similar improvements are observed in other experimental settings.
Fig. 1 presents qualitative comparisons, further illustrating the superiority of DAASH in generating perceptually plausible yet highly effective adversarial examples.
@@
  • We propose DAASH, a fully differentiable, multi-stage framework that learns to combine $\ell_{p}$-norm–constrained base attacks using soft attention.
  • We introduce a novel meta-loss that decouples attack effectiveness and human perceptual quality from the constraints of individual norm-based attacks, enabling better identification of attack strengths across models.
  • We validate DAASH across multiple datasets (CIFAR-10, CIFAR-100, ImageNet), seven robust models, and four post-processing defenses demonstrating substantial improvements over prior perceptual attacks.
@@ 2 Background and Threat Model
Imperceptibility constraints. In the equation above, $d(\cdot,\cdot)$ represents a distance metric and $\epsilon$ defines the perturbation budget. $d(\cdot,\cdot)$ measures the imperceptibility of the perturbation, and the commonly used metrics include $\ell_{2}$ norm Carlini and Wagner (2017) as well as the $\ell_{\infty}$-norm Carlini and Wagner (2017). These $\ell_{p}$-norm constraints may not always align with human perception, and some alternative metrics such as the Structural Similarity Index (SSIM) Wang et al. (2004) is used. SSIM is shown to better capture the luminance, contrast, and structural similarities between two images $\boldsymbol{x}_{1},\boldsymbol{x}_{2}$ and is computed as below:
@@
Besides SSIM, other perceptual similarity metrics that better align with human perception include the Learned Perceptual Image Patch Similarity (LPIPS) Zhang et al. (2018) and the Fréchet Inception Distance (FID) Heusel et al. (2017). In this paper, we use SSIM as the primary metric to optimize for imperceptibility, but we find that using LPIPS and FID yields similar results in practice.
Threat model. In this paper, we focus on the white-box setting, where the attacker is assumed to have complete access to $f(\cdot)$, including the model architecture and parameters. In this setting, untargeted adversarial examples, as defined in Eqn. (1), can be generated via backpropagation using well-defined loss functions with respect to the input $\boldsymbol{x}$. The white-box assumption is widely adopted in the extensive literature Carlini and Wagner (2017); Madry et al. (2018a); Croce and Hein (2020b).
@@ 3 Related Work
Classical approaches for generating adversarial examples include FGSM Goodfellow et al. (2015a), I-FGSM Kurakin et al. (2016), and PGD Madry et al. (2018a), all of which strictly enforce a perturbation budget $\epsilon$ measured in the $\ell_{\infty}$-norm. In contrast, the CW attack Carlini and Wagner (2017) softly enforces an $\ell_{2}$-norm perturbation constraint by incorporating the perturbation magnitude directly into the optimization objective. Closer to our approach are AutoAttack Croce and Hein (2020b) and Composite Attacks Mao et al. (2021). AutoAttack employs an ensemble of a few selected attack methods, while Composite Attacks combine multiple attacks using genetic algorithms. However, these approaches still operate under conventional $\ell_{p}$-norm constraints and rely on simplistic or heuristic strategies for combining base attacks. In contrast, our proposed method DAASH introduces an optimization-based meta-framework to adaptively compose base attacks for improved attack effectiveness and imperceptibility for human eyes.
Recognizing the limitations of norm-constrained attacks in modeling human perception, recent works have shifted toward leveraging perceptual similarity metrics to guide adversarial example generation. For instance, SSAH Luo et al. (2022) operates in the frequency domain to introduce perturbations in less perceptible high-frequency components, while PerC-AL Zhao et al. (2020) optimizes perceptual color distance, though it relies on manual hyperparameter tuning. More recent approaches exploit diffusion models by injecting perturbations into their embedding spaces, thereby influencing the denoising trajectory to produce adversarial outputs. Representative examples include DiffAttack Chen et al. (2024a), AdvAD Li et al. (2024), and DiffPGD Xue et al. (2023). Despite their innovations, these perceptual attacks are often designed as standalone strategies with limited adaptability, making them less effective across diverse target models. As demonstrated in Table 1, such methods frequently underperform in both attack success and imperceptibility. By contrast, DAASH builds on the extensive literature of norm-constrained attacks—known for their compatibility with a wide range of target models—and combines them in a unified framework that enhances the effectiveness and stealth of the resulting adversarial examples.
4 DAASH Attack Design
Next, we describe the high-level motivation behind the design of DAASH and then introduce the details more formally.
Figure 2: An overview of the proposed Differentiable Adversarial Attack Search (DAASH) methodology.
@@ 4.1 Design motivation
Our attack design is motivated by two key observations: (1) Norm-constrained attacks sometimes exhibit complementary strengths and weaknesses Croce and Hein (2020b). So, combining them can yield better performance, especially in under-explored settings of generating adversarial examples beyond $\ell_{p}$-norm constraints; (2) Adversarial example generation is a non-convex optimization process and is prone to getting stuck in poor local optima Madry et al. (2018a).
Regarding Observation 1, the most natural way to combine different attacks is to take a weighted average of the adversarial examples generated by each individual attack. While a linear combination is simple and practical, it introduces two key challenges: (a) How to determine optimal weights for each base attack, and (b) How to ensure that the resulting adversarial example better aligns with human perception rather than just staying within the $\ell_{p}$ constraints.
To address these, we make the combination weights continuous and learnable. We introduce a meta-loss that jointly optimizes for attack effectiveness and perceptual similarity to human vision. This allows the weights to be adjusted through gradient-based optimization, effectively converting norm-constrained base attacks into perceptually aligned attacks—without needing to modify their core mechanisms.
Regarding Observation 2, we mitigate the risk of poor local optima by introducing a multi-stage chaining process. In each stage, we recombine base attacks using newly learned weights and feed the output adversarial example from the previous stage into the next as input. This design allows the attack to iteratively refine the adversarial perturbation. Prior work on transfer-based attacks has shown the benefits of chaining attacks across stages Suya et al. (2020), but typically only use two stages and assign different attacks to each. In contrast, our method allows arbitrary depth and continuously optimized attack combinations across all stages.
4.2 Design Details
where $\boldsymbol{x}_{a}^{j,i}$ denotes the adversarial example generated by the $i$-th base attack at stage $j$. The softmax operation ensures that the final combination remains within the original $\ell_{p}$-norm constraints, as we aim to preserve the structure of the base attacks without modifying them directly. This linear formulation provides a simple and extensible mechanism for combining base attacks. Future work may explore more advanced combination strategies beyond weighted averaging, or treat individual hyperparameters of base attacks, which are fixed in this work, as additional learnable variables.
After the adversarial example generation in each stage, we apply multi-stage chaining as follows: $\boldsymbol{x}_{a}=\mathcal{A}_{N}(\boldsymbol{x}_{a}^{N-1},y)\circ\cdots\circ\mathcal{A}_{1}(\boldsymbol{x}^{0},y)$ where the composition operator $\circ$ indicates the output of each stage is passed as input to the next. This enables iterative refinement of the adversarial perturbation over multiple stages. While we adopt a simple linear chaining strategy here, more sophisticated chaining mechanisms (e.g., cyclic or attention) are possible and left as future work.
Finally, we define the meta-loss that guides the optimization of the attack weights: $\mathcal{L}_{\text{total}}=\lambda_{\text{asr}}\cdot f_{y}(\boldsymbol{x}_{a})+\lambda_{\text{ssim}}\cdot(1-\text{SSIM}(\boldsymbol{x},\boldsymbol{x}_{a}))$ The first term encourages attack effectiveness by maximizing the confidence drop in the ground-truth class $y$ for the adversarial example $\boldsymbol{x}_{a}$. The second term penalizes perceptual distortion, measured using the Structural Similarity Index Measure (SSIM), which ranges from 0 to 1, with higher values indicating greater similarity. In practice, we observe that substituting SSIM with other perceptual similarity metrics yields comparable results. Our full attack algorithm is summarized in Algorithm 1 and a visual demonstration is also shown in Figure 2.
@@ 4.2 Design Details
4:  Sample a mini-batch $\{(\boldsymbol{x},y)\}\subset\mathcal{D}$
@@
Table 1: Performance comparison of adversarial attacks across different datasets and RobustBench models. Base is the ASR without applying any post-processing defenses; Avg. is the average of all the ASR values. SSIM values are scaled to [0,100].
Dataset Target Model Attack ASR (↑) SSIM (↑) LPIPS (↓) FID (↓)

### v2

# DASH: A Meta-Attack Framework for Synthesizing Effective and Stealthy Adversarial Examples

@@ Abstract
Numerous techniques have been proposed for generating adversarial examples under strict $\ell_{p}$-norm constraints. However, such norm-bounded examples often fail to align well with human perception, and only recently have a few methods begun specifically exploring perceptually aligned adversarial examples. Moreover, it remains unclear whether insights from $\ell_{p}$-constrained attacks can be effectively leveraged to improve perceptual efficacy. In this paper, we introduce DASH, a differentiable meta-attack framework that generates effective and perceptually aligned adversarial examples by strategically composing existing $\ell_{p}$-based attack methods. DASH operates in a multi-stage fashion: at each stage, it aggregates candidate adversarial examples from multiple base attacks using learned, adaptive weights and propagates the result to the next stage. A novel meta-loss function guides this process by jointly minimizing misclassification loss and perceptual distortion, enabling the framework to dynamically modulate the contribution of each base attack throughout the stages. We evaluate DASH on adversarially trained robust models across CIFAR-10, CIFAR-100, and ImageNet while considering visual perception metrics (e.g. SSIM, FID, LPIPS) in the perturbation budget (instead of $\ell_{p}$-norm). Despite relying solely on $\ell_{p}$-constrained based methods, DASH significantly outperforms state-of-the-art perceptual attacks such as AdvAD, achieving higher attack success rates (e.g., 20.63% improvement) and superior visual quality, as measured by SSIM, LPIPS, and FID (improvements of $\approx 11$, 0.015, and 5.7, respectively). DASH generalizes well to unseen defenses and different white-box/black-box scenarios, making it a practical and strong baseline for evaluating robustness.
@@ 1 Introduction
Adversarial examples against deep neural networks (DNNs), which add imperceptible perturbations to clean images to induce misclassification (targeted/untargeted) [goodfellow2014explaining], have attracted significant attention from the research community. Various attack methods, including FGSM [goodfellow2015explaining], PGD [madry2018towards], and CW [carlini2017towards], as well as adaptive evaluation frameworks such as AutoAttack [croce2020robustbench], have been developed to benchmark both robust and undefended models. However, the majority of these approaches prioritize imperceptibility by constraining perturbations within a small $\ell_{p}$-norm ball, with $\ell_{2}$ and $\ell_{\infty}$ norms being the most common ones.
Yet, minimizing $\ell_{p}$-norm does not always yield high perceptual similarity to the human visual system [zhang2018unreasonable, laidlaw2020perceptual, sharif2018suitability], despite imperceptibility being a fundamental design goal of stealthy adversarial examples. This disconnect has motivated the adoption of perceptually-aligned similarity metrics such as Structural Similarity Index (SSIM) [wang2004image], Learned Perceptual Image Patch Similarity (LPIPS) [zhang2018unreasonable], and Fréchet Inception Distance (FID) [heusel2017gans] to better assess visual quality [li2024advad, chen2024diffusion]. However, attacks optimized under these perceptual metrics often suffer from reduced success rates; particularly against robust, defended models (see Table 1), highlighting a critical trade-off and leaving significant room for improvement.
Designing effective attacks against a wide range of models, especially hardened ones, remains non-trivial and often requires deep, manual effort, an observation consistently echoed in the literature on $\ell_{p}$-norm attacks [carlini2017towards, athalye2018obfuscated, croce2020reliable]. This complexity intensifies when attempting to simultaneously preserve high perceptual fidelity, thus underscoring the need for systematic and automated strategies that can generalize across models while aligning better with human vision.
Figure 1: Comparative visualization of adversarial perturbations across multiple attack methods (including DASH).
@@
Our key insight is that although individual $\ell_{p}$-norms (e.g., $\ell_{2}$, $\ell_{\infty}$) do not fully capture human perception, they each correlate with different perceptual characteristics (e.g., contrast, texture, edge disruption) [sharif2018suitability, laidlaw2020perceptual]. Therefore, by combining attack strategies optimized under different norms in a principled manner, we may achieve not only improved attack success but also better perceptual quality, particularly when targeting defended models where individual attack methods may fall short. Specifically, we propose DASH (Differential Attack SearcH), a fully differentiable, multi-stage framework that learns to compose multiple $\ell_{p}$-norm–constrained base attacks using soft attention. DASH operates in two stages: (1) it computes a soft attention distribution over adversarial examples generated from individual candidate attacks (e.g., PGD, CW) to weight and compose the final adversarial example; (2) it iteratively refines the perturbations across multiple stages by feeding the output from one stage as the input to the next. This iterative chaining is designed to escape poor local minima inherent to the non-convex landscape of adversarial optimization.
The attention weights are optimized end-to-end through a meta-loss that jointly considers (i) attack success (via minimizing the confidence score of the ground-truth class) and (ii) perceptual similarity (via maximizing SSIM). This allows the framework to dynamically adjust the influence of each candidate attack based on the target model. Unlike existing methods that are rigid in balance between imperceptibility (from human perception) and attack success, DASH learns to trade off these objectives adaptively across different models and datasets.
Empirically, DASH achieves a significant improvement in both attack success rate (ASR) and perceptual fidelity compared to the current state-of-the-art. On robust models trained on CIFAR-100 such as [cui2024decoupled], DASH reaches an average ASR of 99.77% with an SSIM score of 94.43, significantly outperforming state-of-the-art perceptual attacks such as AdvAD [li2024advad] (ASR: 79.14%, SSIM: 83.18) and DiffAttack [chen2024diffusion] (ASR: 77.24%, SSIM: 91.16). Similar improvements are observed in other experimental settings. DASH also transfers well across different models, which demonstrates its effectiveness in black-box settings.
Fig. 1 presents qualitative comparisons, further illustrating the superiority of DASH in generating perceptually plausible yet highly effective adversarial examples.
@@
  • Propose DASH, a differentiable, multi-stage meta-framework that learns to combine multiple base attacks using soft attention. DASH decouples the objectives of attack effectiveness and stealth (e.g., SSIM), allowing for custom optimization of each using a novel meta-loss.
  • Validate DASH across multiple datasets (CIFAR-10, CIFAR-100, ImageNet), seven robust models, and four post-processing defenses demonstrating substantial improvements over prior perceptual attacks.
  • Demonstrate the effectiveness of DASH in black-box settings through rigorous transferability analysis.
@@ 2 Background and Threat Model
Imperceptibility constraints. In the equation above, $d(\cdot,\cdot)$ represents a distance metric and $\epsilon$ defines the perturbation budget. $d(\cdot,\cdot)$ measures the imperceptibility of the perturbation, and the commonly used metrics include $\ell_{2}$ norm [carlini2017towards] as well as the $\ell_{\infty}$-norm [carlini2017towards]. These $\ell_{p}$-norm constraints may not always align with human perception, and some alternative metrics such as the Structural Similarity Index (SSIM) [wang2004image] is used. SSIM is shown to better capture the luminance, contrast, and structural similarities between two images $\boldsymbol{x}_{1},\boldsymbol{x}_{2}$ and is computed as below:
@@
Besides SSIM, other perceptual similarity metrics that better align with human perception include the Learned Perceptual Image Patch Similarity (LPIPS) [zhang2018unreasonable] and the Fréchet Inception Distance (FID) [heusel2017gans]. In this paper, we use SSIM as the perturbation constraint, it’s the primary metric to optimize for imperceptibility, but we find that using LPIPS also yields similar results in practice. Using FID in the optimize process is possible but not ideal in our current experimental setup (small batch FID calculation can be noisy).
Threat model. In this paper, we focus on both white-box and black-box settings. For white-box, the attacker is assumed to have complete access to $f(\cdot)$, including the model architecture and parameters. In this setting, untargeted adversarial examples, as defined in Eqn. (1), can be generated via backpropagation using well-defined loss functions with respect to the input $\boldsymbol{x}$. White-box assumption is the default for robustness analysis (e.g CW [carlini2017towards], AutoAttack [croce2020reliable]) and is also considered in recent (closely related) perception-aligned attacks such as DiffPGD [xue2023diffpgd] (NeurIPS’23) and AdvAD [li2024advad] (NeurIPS’24). For black-box, It is assumed that the attacker does not have any access to the target model. We consider generating the adversarial examples using a surrogate model and then it’s effect is measured on the target model. Note that transfer attacks are under even stricter threat model than query-based black-box attacks.
@@ 3 Related Work
Classical approaches for generating adversarial examples include FGSM [goodfellow2014explaining], I-FGSM [kurakin2016adversarial], and PGD [madry2017towards], all of which strictly enforce a perturbation budget $\epsilon$ measured in the $\ell_{\infty}$-norm. In contrast, the CW attack [carlini2017towards] softly enforces an $\ell_{2}$-norm perturbation constraint by incorporating the perturbation magnitude directly into the optimization objective. Closer to our approach are AutoAttack [croce2020reliable] and Composite Attacks [mao2021composite]. AutoAttack employs an ensemble of a few selected attack methods, while composite attacks combine multiple attacks using genetic algorithms. However, these approaches still operate under conventional $\ell_{p}$-norm constraints and rely on simplistic or heuristic strategies for combining base attacks. In contrast, our proposed method DASH introduces an optimization-based meta-framework to adaptively compose base attacks for improved attack effectiveness and imperceptibility for human eyes.
Recognizing the limitations of norm-constrained attacks in modeling human perception, recent works have shifted toward leveraging perceptual similarity metrics to guide adversarial example generation. For instance, SSAH [luo2022frequency] operates in the frequency domain to introduce perturbations in less perceptible high-frequency components, while PerC-AL [zhao2020towards] optimizes perceptual color distance, though it relies on manual hyperparameter tuning. More recent approaches exploit diffusion models by injecting perturbations into their embedding spaces, thereby influencing the denoising trajectory to produce adversarial outputs. Representative examples include DiffAttack [chen2023diffusion], AdvAD [li2024advad], and DiffPGD [xue2023diffusion]. Despite their innovations, these perceptual attacks are often designed as standalone strategies with limited adaptability, making them less effective across diverse target models. As demonstrated in Table 1, such methods frequently underperform in both attack success and imperceptibility. By contrast, DASH builds on the extensive literature of norm-constrained attacks, known for their compatibility with a wide range of target models and combines them in a unified framework that enhances the effectiveness and stealth of the resulting adversarial examples.
4 DASH Attack Design
Next, we describe the high-level motivation behind the design of DASH and then introduce the details more formally.
Figure 2: An overview of the proposed Differentiable Attack Search (DASH) methodology.
@@ 4.1 Design motivation
Our attack design is motivated by two key observations: (1) Norm-constrained attacks sometimes exhibit complementary strengths and weaknesses [croce2020reliable]. So, combining them can yield better performance, especially in under-explored settings of generating adversarial examples beyond $\ell_{p}$-norm constraints; (2) Adversarial example generation is a non-convex optimization process and is prone to getting stuck in poor local optima [madry2017towards].
Regarding Observation 1, the most natural way to combine different attacks is to take a weighted average of the adversarial examples generated by each individual attack. While a linear combination is practical, it introduces two key challenges: (a) How to determine optimal weights for each base attack, and (b) How to ensure that the resulting adversarial example better aligns with human perception rather than just staying within the $\ell_{p}$ constraints.
To address these, we make the combination weights continuous and learnable. We introduce a meta-loss that jointly optimizes for attack effectiveness and perceptual similarity to human vision. This allows the weights to be adjusted through gradient-based optimization, effectively converting norm-constrained base attacks into perceptually aligned attacks, without needing to modify their core mechanisms.
Regarding Observation 2, we mitigate the risk of poor local optima by introducing a multi-stage chaining process. In each stage, we recombine base attacks using newly learned weights and feed the output adversarial example from the previous stage into the next as input. This design allows the attack to iteratively refine the adversarial perturbation. Prior work on transfer-based attacks has shown the benefits of chaining attacks across stages [suya2020hybrid], but typically only use two stages and assign different attacks to each. In contrast, our method allows arbitrary depth and continuously optimized attack combinations across all stages.
@@
4:  Sample a mini‑batch $\{(\boldsymbol{x},y)\}\subset\mathcal{D}$
@@
4.2 Design Details
where $\boldsymbol{x}_{a}^{j,i}$ denotes the adversarial example generated by the $i$-th base attack at stage $j$. This linear formulation, with softmax probabilities providing normalized importance weights, provides a simple and extensible mechanism for combining base attacks while preserving the overall perceptual similarities. Future work may explore more advanced combination strategies beyond weighted averaging, or treat individual hyperparameters of base attacks, which are fixed in this work, as additional learnable variables.
After the adversarial example generation in each stage, we apply multi-stage chaining as follows: $\boldsymbol{x}_{a}=\mathcal{A}_{N}(\boldsymbol{x}_{a}^{N-1},y)\circ\cdots\circ\mathcal{A}_{1}(\boldsymbol{x}^{0},y)$ where the composition operator $\circ$ indicates the output of each stage is passed as input to the next. This enables iterative refinement of the adversarial perturbation over multiple stages. While we adopt a linear chaining strategy here, other chaining mechanisms (e.g., cyclic or attention) are possible and left as future work.
Finally, we define the meta-loss that guides the optimization of the attack weights: $\mathcal{L}_{\text{total}}=\lambda_{\text{asr}}\cdot f_{y}(\boldsymbol{x}_{a})+\lambda_{\text{ssim}}\cdot(1-\text{SSIM}(\boldsymbol{x},\boldsymbol{x}_{a}))$. The first term encourages attack effectiveness by maximizing the confidence drop in the ground-truth class $y$ for the adversarial example $\boldsymbol{x}_{a}$. The second term penalizes perceptual distortion, measured using the Structural Similarity Index Measure (SSIM), which ranges from 0 to 1, with higher values indicating greater similarity. We include SSIM in the meta-loss to constraint DASH adversarial examples within a visual perturbation budget. While it is not possible to strictly clip an example to a specific SSIM/FID/LPIPS budget limit, the computation of these metrics remains differentiable and hence can be used as an optimization parameter. In the experiment, the weights are optimized until some threshold SSIM is not achieved. In practice, we observe that substituting SSIM with other perceptual similarity metrics yields comparable results. Our full attack algorithm is summarized in Algorithm 1 and a visual demonstration is also shown in Figure 2.
5 Experiments and Results
Next we discuss our core experimental results. All results are obtained using NVIDIA A100 GPU (unless specified). Significant amount of additional experimental results and details (e.g. black-box attacks as DASH base attacks, DASH attack weight distributions, base attack pool variations) can be found inside the supplementary document (appendix). Source code for DASH is also provided as supplementary materials.
5.1 Experimental Setup
We discuss the datasets, robust models, post-processing defenses, metrics, base attacks, and DASH training.
5.1.1 Datasets and Robust Models
Our experiments are conducted on three commonly used benchmark datasets of different scales: CIFAR-10 [krizhevsky2014cifar], CIFAR-100 [Krizhevsky09learningmultiple], and ImageNet [deng2009imagenet]. For each dataset, we randomly sample 1,000 test images. Since standard models without explicit defenses are easily broken by existing attacks, our evaluation focuses on robust models curated by RobustBench [croce2020robustbench], a widely used benchmark for adversarial robustness in which the models are variants of standard adversarial training [madry2017towards].
Table 1: Performance of attacks across datasets and RobustBench [croce2020robustbench] models. Base is the ASR without any post-processing defenses; Avg. is the average of all the ASR values. SSIM values are scaled to [0,100]. Time (sec) is the average generation time of one example.
Dataset Target Model Attack Time (s) ASR (↑) SSIM (↑) LPIPS (↓) FID (↓)